OpenStudy (anonymous):

How do you prove that tan^2 θ•cos^2 θ + cos^2 θ = 1?

OpenStudy (anonymous):

\[\tan^2\theta=\frac{\sin^2\theta}{\cos^2\theta}\]

OpenStudy (anonymous):

Where do I go from there?

OpenStudy (anonymous):

substitute it in the equation

OpenStudy (anonymous):

\[\tan^2\theta\cos^2\theta=\frac{\sin^2\theta}{\cos^2\theta}\times\cos^2\theta=\cdots\] So you have on the left side \[\cdots+\cos^2\theta\]

OpenStudy (anonymous):

So (sin2 θ/cos2 θ)•cos2 θ + cos2 θ = 1

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

How do I simplify it to equal 1? I have to show my work or my teacher won't grade it

OpenStudy (anonymous):

I think I got it. Don't answer as yet @ranga, I might need you soon

OpenStudy (ranga):

I will use x instead of theta as it is easier to type. tan^2 x•cos^2 x + cos^2 x = sin^2(x) / cos^2(x) * cos^2(x) + cos^2(x) = sin^2(x) + cos^2(x) There is a trigonometric identity that says: sin^2(x) + cos^2(x) = 1

OpenStudy (anonymous):

Ok well I guess I didn't have it lol. My answer looked like sin2 θ/cos2 θ=1

OpenStudy (anonymous):

Thank you once more @ranga. You're awesome

OpenStudy (ranga):

You are welcome.